Fermi National Laboratory
 

Accelerator Update for March 11, 2009 - March 13, 2009

 

Wednesday March 11

The day shift began with Operations monitoring the Tevatron (TeV) store 6888 (117.17E30) and the Recycler stash (277.4E10), with the Antiproton Source stacking (16.9E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

A Booster expert began tuning and investigating bad beam positions at 8:40 AM.

At 9:06 AM, two MI kickers (I:KPS4A and B) tripped off. They reset.

At 9:24 AM, Operations reported that a Linac quadrupole power supply (L:QPS117) had made two setting jumps. An operator made adjustments for each of these jumps.

At 10:14 AM, Operations reported more problems with L:QPS117. It was adjusted back to its normal setting.

At 10:18 AM, a MiniBooNE vacuum pump (IP8667) tripped off. It reset.

TeV experts began an end-of-store study around 11:02 AM.

At 11:16 AM, the P1 transfer line beam permit tripped off due to a power supply trip (V714). Everything reset.

Operations put the TeV into shot setup at 12:16 PM.

Linac experts began working on L:QPS117 at 1:13 PM.

Operations established store 6889 at 1:39 PM with an initial luminosity of 310.97E30.

Pbar power supply experts began working on A:IB and HVF12 at 2:06 PM. This put a halt to stacking until 2:44 PM when they completed their work.

At 3:06 PM, Operations reported that when Pbar returned to stacking mode, the NuMI beam positions were off. Operations halted stacking and contacted an expert.

Pbar experts reported at 3:37 PM that something in the timeline had caused problems with the mixed mode autotune for NuMI.

Pbar resumed stacking at 3:45 PM.

At 3:49 PM, L:QPS114 tripped off. It reset.

Operations reported at 4:03 PM that the Recycler had stashed antiprotons five times during the day shift.

At 7:47 PM, two Booster vacuum pumps (IPRF17 and 171) tripped off and caused two RF stations (BRF7 and 8) to also trip. Everything reset.

At 10:24 PM, BRF16 was tripping off repeatedly. Operators bypassed the station and brought BRF14 on its place.

Operations reported at 11:59 PM that the Recycler had stashed eight times during the evening shift.

Thursday March 12

The midnight shift began with Operations monitoring store 6889 (99.34E30) and the stash (249.9E10), with Pbar stacking (28.1E10), and with NuMI and MiniBooNE taking beam.

At 12:20 AM, two Booster vacuum pumps tripped off (IPRF61 and 72). They reset.

At 1:31 AM, a Linac quadrupole power supply (QPS102) tripped off. It reset.

Operations terminated store 6889 at 4:58 AM.

Operations put the TeV into shot setup at 5:10 AM.

At 6:59 AM, a Linac Klystron RF station (KRF3) tripped off. It reset.

Operations established store 6890 at 7:06 AM with an initial luminosity of 308.27E30.

Operations reported at 7:54 AM that the Recycler had stashed six times during the midnight shift.

Switchyard (SY) experts began tuning beam to the SY dump at 8:52 AM.

At 10:27 AM, a Linac quadrupole power supply (L:QPS226) was drifting high. Operators adjusted it.

Operations reported at 11:10 AM that the Meson MTest T988 “AIRFLY” experiment had been approved to take beam. Operations began sending them beam at 11:15 AM.

At 1:32 PM, Linac experts began working on L:QPS102 and LRF2 and 4. Preaccelerator experts began working on the I- Source PLC. Operations turned off Linac for this work.

A Booster expert began taking Booster chopper measurements at 1:39 PM.

A Pbar expert began checking for a possible LCW leak in and around the water cage around 2:09 PM. He was out about an hour later. He did not find a leak.

Preaccelerator experts completed their work at 2:24 PM.

At 2:36 PM, BRF3 tripped off. An expert investigated.

Linac resumed sending beam at 3:31 PM.

Operations reported at 4:06 PM that the Recycler had stashed five times during the day shift.

At 5:29 PM, Operations reported that not many antiprotons were extracted out of the Accumulator during the first transfer of a stash to the Recycler. On the second try, antiprotons made it to the extraction orbit, but never made it to MI. Operations halted the stash to investigate and contacted an expert. Pbar resumed stacking.

Operations disabled the MTest beam at 6:15 PM.

Operations reported at 6:33 PM that they had successfully stashed antiprotons.

Operations terminated store 6890 at 11:21 PM.

Operations put the TeV into shot setup at 11:30 PM.

At 11:50 PM, Operations reported that there were some devices in the TeV alarm list showing errors. This had caused problems for the TeV sequencer.

Operations reported at 11:51 PM that the Recycler had stashed eight times during the evening shift.

Friday March 13

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (236.8E10), with Pbar stacking (51.5E10), and with NuMI and MiniBooNE taking beam.

Operations established store 6891 at 1:18 AM with an initial luminosity of 298.94E30.

At 7 AM, the 400 MeV area beam permit tripped off for the first of three times in 13 minutes due to bad beam positions. It reset.

Operations reported at 7:57 AM that the Recycler had stashed six times during the midnight shift.

The Plans for Friday and the Weekend

At 8:20 AM, a D0 transformer trip caused a power glitch. Store 6891 was lost. The TeV suffered a quench at sector C4. FESS reported that they would have to shut off power to TeV sectors D0, D1, D2, and D3 to make repairs. The plan for today is to bring power back to D0, repair any associated damages, and then return to stacking, stashing, and storing. The Run coordinator will allow accelerator maintenance during the D0 repairs.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/13/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory